OpenStudy (anonymous):

if $7000 is invested at a rate of 6% compounded continuously, find the balance in the account after 7 years

OpenStudy (anonymous):

this is a simple problem just google the formula for compound interest and substitute your values.

OpenStudy (anonymous):

a= p(1+r/t)^nt

OpenStudy (anonymous):

but i dont know how to plug it in like you said

OpenStudy (anonymous):

a= p(1+r/t)^nt ,this for compouded semiannually for compounded continuously you use a=7000.e^(0.06*7)

OpenStudy (anonymous):

I hpoe it help

OpenStudy (anonymous):

for compouded semiannually you plug a=7000(1+0.06/2)^(2*7) for compounded monthly you use this; a= 7000(1+0.06/12)^(12*7)